# A Compute and Communication Runtime Model for Loihi 2

### v1

@@ Abstract
Neuromorphic computers hold the potential to vastly improve the speed and efficiency of a wide range of computational kernels with their asynchronous, compute-memory co-located, spatially distributed, and scalable nature. However, performance models that are simple yet sufficiently expressive to predict runtime on actual neuromorphic hardware are lacking, posing a challenge for researchers and developers who strive to design fast algorithms and kernels. As breaking the memory bandwidth wall of conventional von-Neumann architectures is a primary neuromorphic advantage, modeling communication time is especially important. At the same time, modeling communication time is difficult, as complex congestion patterns arise in a heavily-loaded Network-on-Chip. In this work, we introduce the first max-affine lower-bound runtime model—a multi-dimensional roofline model—for Intel’s Loihi 2 neuromorphic chip that quantitatively accounts for both compute and communication based on a suite of microbenchmarks. Despite being a lower-bound model, we observe a tight correspondence (Pearson correlation coefficient greater than or equal to 0.97) between our model’s estimated runtime and the measured runtime on Loihi 2 for a neural network linear layer, i.e., matrix-vector multiplication, and for an example application, a Quadratic Unconstrained Binary Optimization solver. Furthermore, we derive analytical expressions for communication-bottlenecked runtime to study scalability of the linear layer, revealing an area-runtime tradeoff for different spatial workload configurations with linear to superliner runtime scaling in layer size with a variety of constant factors. Our max-affine runtime model helps empower the design of high-speed algorithms and kernels for Loihi 2.
@@ I Introduction
However, high-level benchmarks entangle hardware and algorithm innovation, which can impede hardware experts and algorithm experts from independently advancing their respective areas of expertise. Well-designed abstractions facilitate harmonious development across technology stack layers—from software, AI frameworks, hardware, compilers, and so on. Indeed, these interfacing abstractions are routinely credited with accelerating development [10, 14]. In neuromorphic computing, while there certainly will always be an irreducible interdisciplinary component of algorithm-hardware co-design, the lack of a precise quantitative abstraction through which hardware and algorithms experts can readily exchange and build upon each others’ innovations suggests that a layer of abstraction is missing. For algorithm-hardware co-design, the missing layer of abstraction is a performance model, i.e., a simplified model of the hardware that quantitatively predicts the performance of an algorithm (runtime, energy, etc.). Indeed, such performance models are the workhorse of kernel developers and hardware designers in mainstream architectures. For example, the roofline model for GPUs, which distinguishes between compute-bound and memory-bandwidth-bound kernels and provides a quantitative lower-bound on kernel runtime, is routinely used by kernel writers to optimize performance and by hardware designers to understand the impact of design tradeoffs [23, 24].
@@
We find that, although our max-affine model produces a lower-bound on runtime, its predicted runtime is near and highly correlated with measured runtime on Loihi 2. This holds for a variety of linear layer shapes, sizes, and partitions, including those bottlenecked by compute, communication, or other factors. Importantly, we also find that our model’s predicted runtime is highly correlated with measured runtime for sparse linear layers that stress the communication mesh on Loihi 2, across a variety of spatial placement patterns, further validating the more complex traffic resolution time term in the model. We also demonstrate that our model is predictive of runtime for an example application, a Quadratic Unconstrained Binary Optimization (QUBO) solver [16]. Across all workloads measured in this paper, the Pearson correlation coefficient between measured and estimated runtime is $\geq$ 0.97.
@@ II-1 Loihi 2 Systems
A single Loihi 2 chip is primarily comprised of NeuroCores, Routers, and mesh links (Figure 1). NeuroCores each have memory and compute: they store neuron states and synaptic weights, and they also perform computations, such as neuron updates and synaptic MAC operations. Routers facilitate communication between NeuroCores. Via two bidirectional mesh links, each router is connected to four NeuroCores, and routers are connected to nearest-neighbor routers in the four cardinal directions. Messages from NeuroCores are routed to their destinations according to dimension-order (X-Y) routing, as illustrated in Figure 1. The NoC also supports a synchronization operation, known as barrier synchronization, to synchronize algorithmic timesteps across NeuroCores: for many neuromorphic algorithms, e.g., a leak-integrate-and-fire network [21] evolving in discretized time, execution proceeds in a sequence of algorithmic timesteps, and it is critical that NeuroCores are synchronized to this degree. Other cores, such as CPU cores, are used for management or other special functions. NeuroCores, routers, and mesh links handle the brunt of Loihi 2 workloads, so they are the focus of our runtime model.
@@ II-2 Loihi 2 Profiling Tools
Communication, however, cannot be inferred from simple hardware counters, as messages travel via various paths. We therefore introduce a tool called TrafficStats which infers the bandwidth used on every mesh link. TrafficStats inspects the configured NeuroCores and their message programs to infer how much message bandwidth will be loaded on every mesh link, taking into account dimension-order routing for each message. TrafficStats assumes a uniform message-send frequency, i.e., that every neuron spikes on every algorithmic time-step, which is the case for the linear layer workloads we consider in this work. However, TrafficStats can also be used to compute mesh link bandwidth loads for more specific activity patterns. E.g., we apply an activation sparsity factor to estimate traffic for the QUBO solver. The TrafficStats tool is available in NxKernel, a flexible, low-level, and performant software package for mapping networks to Loihi 2. Mesh link loads inferred from TrafficStats are visualized in the Loihi 2 traffic schematics in this paper (Figure 3).
@@ III Microbenchmark suite
In order to create a runtime model that is high-level enough to interface with algorithm designers and low-level enough to be predictive of actual hardware performance, we characterize the key compute and communication runtime bottlenecks on Loihi 2 with a suite of microbenchmarks. Our microbenchmark suite specified in NxKernel measures the effective NeuroCore barrier synchornization time, the effective NeuroCore peak operation rate for the chief compute operations SynOps, SynMem reads, and DendOps, and the effective peak bandwidth for NoC mesh links. Measuring effective rates without delving into circuit-level details allows our microbenchmark suite to yield an easily understood and useful characterization of compute and communication.
@@ V Loihi 2 Results
For several workloads and configurations, we measure time per step on Loihi 2 and compare it to our max-affine model’s predicted time per step. We find that the max-affine model indeed reflects a lower-bound estimate for runtime, and moreover, can be near an exact runtime predictor. This close correspondence holds across compute and communication bottlenecks. Furthermore, with a communication-bottlenecked workload, we explore Loihi 2 spatial placement options, which significantly affect mesh traffic, and show that our max-affine model’s predictive power holds across this space.
@@ V-1 Dense Linear Layer
Figure 6: Correspondence between modeled and measured performance for a traffic-dominated tiled identity workload. For the same traffic-dominated workload mapped to the same number of NeuroCores, different placement patterns (symbols/colors) of the NeuroCores vastly alters the runtime as different placement patterns yield different levels of NoC traffic congestion. Over all placement patterns, however, the max-affine model maintains predictive power; the Pearson correlation coefficient is 0.999. †Workloads characterized on an Oheo Gulch system with a N3C2-revision Loihi 2 chip running an unreleased patch of NxCore 2.5.10 and NxKernel 0.4.0.
@@ V-2 Traffic-bottlenecked Linear Layer
To evaluate the runtime model under conditions where communication time dominates and spatial NeuroCore placement choices become important, we run a linear layer workload with a sparse tiled-identity synaptic weight matrix under a variety of placement configurations. We again observe a correspondence between the max-affine model estimated step time and measured step time on Loihi 2; see Figure 6. The tiled identity workload has 8 origin cores sending to 8 destination cores with 1024 neurons per core. The connectivity matrix is $8192\times 8192$, comprised of $1024\times 1024$ identity matrix tiles connecting each origin core’s $1024$ neurons to all neurons in a destination core. The 88 $1024\times 1024$ identity matrix tiles stored on each destination core are sparse-packed, taking little memory. This sparse-packed tiled-identity connectivity ensures that all neurons on all origin cores are connected to all neurons on all destination cores, maximizing Noc traffic, while keeping SynOps and SynMem reads at a minimum. We run this workload under several different placement patterns, including random placement and an X-shaped placement; see Figure 3bc for examples. Crucially, the placement patterns affect the traffic routes from the origin cores to the destination cores—some patterns cause more congestion than others—and this is reflected in our max-affine performance model (Equation 1) as a change in the heaviest link load, $N_{LL}$. The X-shaped pattern exhibited the lowest time per step.
@@ V-3 QUBO Solver
We further demonstrate the predictive power of our max-affine performance model by estimating the runtime of a Quadratic Unconstrained Binary Optimization (QUBO) combinatorial solver [16, 20] (Figure 7). In the QUBO solver, a QUBO problem $\min_{x}x^{T}Qx$ where $x\in{\{0,1\}}^{N}$ is encoded in the states and recurrent connectivity of a population of $N$ neurons; the $i$th neuron encodes $x_{i}$. An additional temperature neuron and loss neuron are connected to this population to facilitate annealing and progress tracking, but the bulk of compute and communication are performed by the population of $N$ neurons. As the QUBO solver runs, each of the $N$ neurons alternate between two stages: checking and switching. In the checking stage, the neurons perform a calculation and coordinate to determine an appropriate probability with which to flip their state variable $x_{i}$ (from 0 to 1, or vice versa). In the switching stage, the neurons stochastically flip their state variable $x_{i}$ according to the previously computed probability and communicate updates to other neurons so that the checking stage can be repeated with an up-to-date state of the QUBO problem. Alternating checking and switching stages continue until the QUBO loss converges to a satisfactory level.
The QUBO solver is a traffic-bottlenecked workload, and we can analyze the checking stage and switching stage separately to form an estimate of the mean time per step in each stage under a variety of QUBO problem sizes and NeuroCore placement patterns; see Figure 7. Importantly, the traffic pattern is the same for both the checking stage and the switching stage, as they use the same recurrent connectivity structure of the QUBO matrix $Q$. Thus, we can use the TrafficStats tool on the $N$ neurons and their recurrent connectivity to determine the heaviest link load. However, importantly, the TrafficStats tool by default assumes that all neurons emit spikes every timestep, but in the QUBO solver, neurons often do not spike, as they spike stochastically. We account for this by measuring the activity rate (1activity sparsity)(1-$\text${activity sparsity}) in the checking stage and switching stage, and multiplying the activity rate by the heaviest link load from TrafficStats. This yields an estimate which corresponds to measured mean time per step for the checking stage and switching stage over a variety of QUBO problem sizes (Figure 7).
@@ VI Analytical model for NoC congestion
We can gain further insight into how different placement patterns of a linear layer workload congest the mesh to different degrees by deriving analytical expressions for the heaviest link load given a placement pattern. For simplicity, we conceptually aggregate the two-mesh bidirectional links on Loihi 2 into effective single-mesh bidirectional links as traffic is typically distributed across both meshes, and we focus on variations of the highest-performing placement pattern, the X-shaped placement. Namely, we consider placement patterns in which each pair of origin and destination cores are assigned to a router and all routers have at most two populated NeuroCores, like in Figure 3c. Intuitively, assigning an origin core and destination core to the same router keeps at least one communication path off the router-to-router mesh links (the communication path between the origin core and the destination core on the same router) and also minimizes any single router’s burden for the total traffic coming into each destination core (each destination core is a sink, and having more than one destination core on a single router doubles the traffic incoming to that router). Importantly, we will see that the minimum possible heaviest link load is achievable under these conditions.
@@
We calculate the heaviest leftward router-to-router link load by deriving an expression for the leftward link load at any router taking in account dimension-order routing of messages from all origin cores. Let us first consider the routing of messages from a populated router ($P_{ij}=1$) at coordinates $(i,j)$; note that we consider $i=1$ the top row of routers and $j=1$ the leftmost column of routers, thus leftward-directed traffic is from higher values of $j$ to lower values of $j$. This router has one origin core, which emits $M$ messages. Each message is destined for a different destination core and its associated router. Because messages are routed horizontally first, we know that the number of messages that route horizontally and then route vertically up and down column $l$ is $\text{colsum}_{l}=\sum_{k=1}^{n}P_{kl}$, i.e., all messages destined for a given column must arrive in that column via horizontal routing. Now, consider the leftmost router in a row $i$: $(i,1)$. There is no leftward traffic as it is on the edge of the NoC. Next, consider the second leftmost router in row $i$: $(i,2)$. $\text{colsum}_{1}$ remains to be routed vertically at the leftmost router $(i,1)$, so the leftward traffic from router $(i,2)$ is $\text{colsum}_{1}$. Next, consider the third leftmost router in row $i$: $(i,3)$. $\text{colsum}_{1}+\text{colsum}_{2}$ remains to be routed either vertically at $(i,2)$, or $(i,1)$, thus the leftward traffic from router $(i,3)$ is $\text{colsum}_{1}+\text{colsum}_{2}$. This is a cumulative sum, and this pattern continues until $l=j$; at $l=j+1$, traffic is now traveling to the right as the origin router is in column $j$, and thus leftward traffic at node $(i,j+1)$ is zero, and similarly for all coordinates further right. We write this cumulative sum for the leftward traffic at $(i,l)$ from the router $(i,j)$, take a sum over all $j$ to get the total leftward traffic from all routers in row $i$, and take the max over all leftward links $(i,l)$ to obtain the heaviest leftward router-to-router link load:
@@
With Equation 2, we can plug in particular placement patterns and explore how the heaviest router-to-router link load grows with the hypothetical size of the NoC. For a saturated rectangle, $P_{ij}=1\;\forall 1\leq i\leq n,1\leq j\leq m$, the heaviest leftward link load is $N_{LL}=nm^{2}/4$. A similar calculation for upward link load yields a symmetric formula in $n$ and $m$, and this implies the intuitive result that the optimal rectangular placement is a square with $n=m$; $n\neq m$ leaves room to reshape the rectangle to achieve a lower heaviest link load. For the square, the number of origin-destination core pairs is $M=\sum_{i,j}P_{ij}=n^{2}$, and so the heaviest link load is $\frac{M^{3/2}}{4}$. In contrast, an X-shaped placement $P_{ij}=\delta_{ij}+\delta_{i,n-j+1}\,\forall i,j\in{0,...,n}$, where $n$ is evenly divisible by $2$, yields $N_{LL}=2(n-1)$. In the X-shaped placement, $M=2n$, and thus the heaviest router-to-router link load is M2M-2. As noted earlier, the core-to-router link load is already $M$ for all placement patterns, thus the X-shape’s M2M-2 heaviest router-to-router link load achieves the minimum possible heaviest link load overall.
@@
Figure 8: Heaviest router-to-router link load and NoC area as the number of processors scales for different placement patterns. A tradeoff exists in placement pattern compactness and speed: runtime grows linearly and area grows superlinearly for the X-shaped placement pattern, and vice versa for the satruated square placement pattern.

### v2

# arXiv:2601.10035v2[cs.NE] 27 Feb 2026

@@ Abstract
Neuromorphic computers hold the potential to vastly improve the speed and efficiency of a wide range of computational kernels with their asynchronous, compute-memory co-located, spatially distributed, and scalable nature. However, performance models that are simple yet sufficiently expressive to predict runtime on actual neuromorphic hardware are lacking, posing a challenge for researchers and developers who strive to design fast algorithms and kernels. As breaking the memory bandwidth wall of conventional von-Neumann architectures is a primary neuromorphic advantage, modeling communication time is especially important. At the same time, modeling communication time is difficult, as complex congestion patterns arise in a heavily-loaded Network-on-Chip. In this work, we introduce the first max-affine lower-bound runtime model—a multi-dimensional roofline model—for Intel’s Loihi 2 neuromorphic chip that quantitatively accounts for both compute and communication based on a suite of microbenchmarks. Despite being a lower-bound model, we observe a tight correspondence (Pearson correlation coefficient greater than or equal to 0.97) between our model’s estimated runtime and the measured runtime on Loihi 2 for a neural network linear layer, i.e., matrix-vector multiplication, and for an example application, a Quadratic Unconstrained Binary Optimization solver. Furthermore, we derive analytical expressions for communication-bottlenecked runtime to study scalability of the linear layer, revealing an area-runtime tradeoff for different spatial workload configurations with linear to superlinear runtime scaling in layer size with a variety of constant factors. Our max-affine runtime model helps empower the design of high-speed algorithms and kernels for Loihi 2.
@@ I Introduction
However, high-level benchmarks entangle hardware and algorithm innovation, which can impede hardware experts and algorithm experts from independently advancing their respective areas of expertise. Well-designed abstractions facilitate harmonious development across technology stack layers—from software, AI frameworks, hardware, compilers, and so on. Indeed, these interfacing abstractions are routinely credited with accelerating development [10, 14]. In neuromorphic computing, while there certainly will always be an irreducible interdisciplinary component of algorithm-hardware co-design, the lack of a precise quantitative abstraction through which hardware and algorithms experts can readily exchange and build upon each other’s innovations suggests that a layer of abstraction is missing. For algorithm-hardware co-design, the missing layer of abstraction is a performance model, i.e., a simplified model of the hardware that quantitatively predicts the performance of an algorithm (runtime, energy, etc.). Indeed, such performance models are the workhorse of kernel developers and hardware designers in mainstream architectures. For example, the roofline model for GPUs, which distinguishes between compute-bound and memory-bandwidth-bound kernels and provides a quantitative lower-bound on kernel runtime, is routinely used by kernel writers to optimize performance and by hardware designers to understand the impact of design tradeoffs [23, 24].
@@
We find that, although our max-affine model produces a lower-bound on runtime, its predicted runtime is close to and highly correlated with measured runtime on Loihi 2. This holds for a variety of linear layer shapes, sizes, and partitions, including those bottlenecked by compute, communication, or other factors. Importantly, we also find that our model’s predicted runtime is highly correlated with measured runtime for sparse linear layers that stress the communication mesh on Loihi 2, across a variety of spatial placement patterns, further validating the more complex traffic resolution time term in the model. We also demonstrate that our model is predictive of runtime for an example application, a Quadratic Unconstrained Binary Optimization (QUBO) solver [16]. Across all workloads measured in this paper, the Pearson correlation coefficient between measured and estimated runtime is $\geq$ 0.97.
@@ II-1 Loihi 2 Systems
A single Loihi 2 chip is primarily composed of NeuroCores, Routers, and mesh links (Figure 1). NeuroCores each have memory and compute: they store neuron states and synaptic weights, and they also perform computations, such as neuron updates and synaptic MAC operations. Routers facilitate communication between NeuroCores. Via two bidirectional mesh links, each router is connected to four NeuroCores, and routers are connected to nearest-neighbor routers in the four cardinal directions. Messages from NeuroCores are routed to their destinations according to dimension-order (X-Y) routing, as illustrated in Figure 1. The NoC also supports a synchronization operation, known as barrier synchronization, to synchronize algorithmic timesteps across NeuroCores: for many neuromorphic algorithms, e.g., a leak-integrate-and-fire network [21] evolving in discretized time, execution proceeds in a sequence of algorithmic timesteps, and it is critical that NeuroCores are synchronized to this degree. Other cores, such as CPU cores, are used for management or other special functions. NeuroCores, routers, and mesh links handle the brunt of Loihi 2 workloads, so they are the focus of our runtime model.
@@ II-2 Loihi 2 Profiling Tools
Communication, however, cannot be inferred from simple hardware counters, as messages travel via various paths. We therefore introduce a tool called TrafficStats which infers the bandwidth used on every mesh link. TrafficStats inspects the configured NeuroCores and their message programs to infer how much message bandwidth will be loaded on every mesh link, taking into account dimension-order routing for each message. TrafficStats assumes a uniform message-send frequency, i.e., that every neuron spikes on every algorithmic timestep, which is the case for the linear layer workloads we consider in this work. However, TrafficStats can also be used to compute mesh link bandwidth loads for more specific activity patterns. E.g., we apply an activation sparsity factor to estimate traffic for the QUBO solver. The TrafficStats tool is available in NxKernel, a flexible, low-level, and performant software package for mapping networks to Loihi 2. Mesh link loads inferred from TrafficStats are visualized in the Loihi 2 traffic schematics in this paper (Figure 3).
@@ III Microbenchmark suite
In order to create a runtime model that is high-level enough to interface with algorithm designers and low-level enough to be predictive of actual hardware performance, we characterize the key compute and communication runtime bottlenecks on Loihi 2 with a suite of microbenchmarks. Our microbenchmark suite specified in NxKernel measures the effective NeuroCore barrier synchronization time, the effective NeuroCore peak operation rate for the chief compute operations SynOps, SynMem reads, and DendOps, and the effective peak bandwidth for NoC mesh links. Measuring effective rates without delving into circuit-level details allows our microbenchmark suite to yield an easily understood and useful characterization of compute and communication.
@@ V Loihi 2 Results
For several workloads and configurations, we measure time per step on Loihi 2 and compare it to our max-affine model’s predicted time per step. We find that the max-affine model indeed reflects a lower-bound estimate for runtime, and moreover, can be a near-exact runtime predictor. This close correspondence holds across compute and communication bottlenecks. Furthermore, with a communication-bottlenecked workload, we explore Loihi 2 spatial placement options, which significantly affect mesh traffic, and show that our max-affine model’s predictive power holds across this space.
@@ V-1 Dense Linear Layer
Figure 6: Correspondence between modeled and measured performance for a traffic-dominated tiled identity workload. For the same traffic-dominated workload mapped to the same number of NeuroCores, different placement patterns (symbols/colors) of the NeuroCores vastly alter the runtime as different placement patterns yield different levels of NoC traffic congestion. Over all placement patterns, however, the max-affine model maintains predictive power; the Pearson correlation coefficient is 0.999. †Workloads characterized on an Oheo Gulch system with a N3C2-revision Loihi 2 chip running an unreleased patch of NxCore 2.5.10 and NxKernel 0.4.0.
@@ V-2 Traffic-bottlenecked Linear Layer
To evaluate the runtime model under conditions where communication time dominates and spatial NeuroCore placement choices become important, we run a linear layer workload with a sparse tiled-identity synaptic weight matrix under a variety of placement configurations. We again observe a correspondence between the max-affine model estimated step time and measured step time on Loihi 2; see Figure 6. The tiled identity workload has 8 origin cores sending to 8 destination cores with 1024 neurons per core. The connectivity matrix is $8192\times 8192$, composed of $1024\times 1024$ identity matrix tiles connecting each origin core’s $1024$ neurons to all neurons in a destination core. The 88 $1024\times 1024$ identity matrix tiles stored on each destination core are sparse-packed, taking little memory. This sparse-packed tiled-identity connectivity ensures that all neurons on all origin cores are connected to all neurons on all destination cores, maximizing NoC traffic, while keeping SynOps and SynMem reads at a minimum. We run this workload under several different placement patterns, including random placement and an X-shaped placement; see Figure 3bc for examples. Crucially, the placement patterns affect the traffic routes from the origin cores to the destination cores—some patterns cause more congestion than others—and this is reflected in our max-affine performance model (Equation 1) as a change in the heaviest link load, $N_{LL}$. The X-shaped pattern exhibited the lowest time per step.
@@ V-3 QUBO Solver
We further demonstrate the predictive power of our max-affine performance model by estimating the runtime of a Quadratic Unconstrained Binary Optimization (QUBO) combinatorial solver [16, 20] (Figure 7). In the QUBO solver, a QUBO problem $\min_{x}x^{T}Qx$ where $x\in{\{0,1\}}^{N}$ is encoded in the states and recurrent connectivity of a population of $N$ neurons; the $i$th neuron encodes $x_{i}$. An additional temperature neuron and loss neuron are connected to this population to facilitate annealing and progress tracking, but the bulk of compute and communication are performed by the population of $N$ neurons. As the QUBO solver runs, each of the $N$ neurons alternates between two stages: checking and switching. In the checking stage, the neurons perform a calculation and coordinate to determine an appropriate probability with which to flip their state variable $x_{i}$ (from 0 to 1, or vice versa). In the switching stage, the neurons stochastically flip their state variable $x_{i}$ according to the previously computed probability and communicate updates to other neurons so that the checking stage can be repeated with an up-to-date state of the QUBO problem. Alternating checking and switching stages continue until the QUBO loss converges to a satisfactory level.
The QUBO solver is a traffic-bottlenecked workload, and we can analyze the checking stage and switching stage separately to form an estimate of the mean time per step in each stage under a variety of QUBO problem sizes and NeuroCore placement patterns; see Figure 7. Importantly, the traffic pattern is the same for both the checking stage and the switching stage, as they use the same recurrent connectivity structure of the QUBO matrix $Q$. Thus, we can use the TrafficStats tool on the $N$ neurons and their recurrent connectivity to determine the heaviest link load. However, importantly, the TrafficStats tool by default assumes that all neurons emit spikes every timestep, but in the QUBO solver, neurons often do not spike, as they spike stochastically. We account for this by measuring the activity rate (1activity sparsity)(1-$\text${activity sparsity}) in the checking stage and switching stage, and multiplying the activity rate by the heaviest link load from TrafficStats. This yields an estimate that corresponds to measured mean time per step for the checking stage and switching stage over a variety of QUBO problem sizes (Figure 7).
@@ VI Analytical model for NoC congestion
We can gain further insight into how different placement patterns of a linear layer workload congest the mesh to different degrees by deriving analytical expressions for the heaviest link load given a placement pattern. For simplicity, we conceptually aggregate the two-mesh bidirectional links on Loihi 2 into effective single-mesh bidirectional links as traffic is typically distributed across both meshes, and we focus on variations of the highest-performing placement pattern, the X-shaped placement. Namely, we consider placement patterns in which each pair of origin and destination cores is assigned to a router and all routers have at most two populated NeuroCores, like in Figure 3c. Intuitively, assigning an origin core and destination core to the same router keeps at least one communication path off the router-to-router mesh links (the communication path between the origin core and the destination core on the same router) and also minimizes any single router’s burden for the total traffic coming into each destination core (each destination core is a sink, and having more than one destination core on a single router doubles the traffic incoming to that router). Importantly, we will see that the minimum possible heaviest link load is achievable under these conditions.
@@
We calculate the heaviest leftward router-to-router link load by deriving an expression for the leftward link load at any router taking into account dimension-order routing of messages from all origin cores. Let us first consider the routing of messages from a populated router ($P_{ij}=1$) at coordinates $(i,j)$; note that we consider $i=1$ the top row of routers and $j=1$ the leftmost column of routers, thus leftward-directed traffic is from higher values of $j$ to lower values of $j$. This router has one origin core, which emits $M$ messages. Each message is destined for a different destination core and its associated router. Because messages are routed horizontally first, we know that the number of messages that route horizontally and then route vertically up and down column $l$ is $\text{colsum}_{l}=\sum_{k=1}^{n}P_{kl}$, i.e., all messages destined for a given column must arrive in that column via horizontal routing. Now, consider the leftmost router in a row $i$: $(i,1)$. There is no leftward traffic as it is on the edge of the NoC. Next, consider the second leftmost router in row $i$: $(i,2)$. $\text{colsum}_{1}$ remains to be routed vertically at the leftmost router $(i,1)$, so the leftward traffic from router $(i,2)$ is $\text{colsum}_{1}$. Next, consider the third leftmost router in row $i$: $(i,3)$. $\text{colsum}_{1}+\text{colsum}_{2}$ remains to be routed either vertically at $(i,2)$, or $(i,1)$, thus the leftward traffic from router $(i,3)$ is $\text{colsum}_{1}+\text{colsum}_{2}$. This is a cumulative sum, and this pattern continues until $l=j$; at $l=j+1$, traffic is now traveling to the right as the origin router is in column $j$, and thus leftward traffic at node $(i,j+1)$ is zero, and similarly for all coordinates further right. We write this cumulative sum for the leftward traffic at $(i,l)$ from the router $(i,j)$, take a sum over all $j$ to get the total leftward traffic from all routers in row $i$, and take the max over all leftward links $(i,l)$ to obtain the heaviest leftward router-to-router link load:
@@
With Equation 2, we can plug in particular placement patterns and explore how the heaviest router-to-router link load grows with the hypothetical size of the NoC. For a saturated rectangle, $P_{ij}=1\;\forall 1\leq i\leq n,1\leq j\leq m$, the heaviest leftward link load is $N_{LL}=nm^{2}/4$. A similar calculation for upward link load yields a symmetric formula in $n$ and $m$, and this implies the intuitive result that the optimal rectangular placement is a square with $n=m$; $n\neq m$ leaves room to reshape the rectangle to achieve a lower heaviest link load. For the square, the number of origin-destination core pairs is $M=\sum_{i,j}P_{ij}=n^{2}$, and so the heaviest link load is $\frac{M^{3/2}}{4}$. In contrast, an X-shaped placement $P_{ij}=\delta_{ij}+\delta_{i,n-j+1}\,\forall i,j\in\{1,...,n\}$, where $n$ is evenly divisible by $2$, yields $N_{LL}=2(n-1)$. In the X-shaped placement, $M=2n$, and thus the heaviest router-to-router link load is M2M-2. As noted earlier, the core-to-router link load is already $M$ for all placement patterns, thus the X-shape’s M2M-2 heaviest router-to-router link load achieves the minimum possible heaviest link load overall.
@@
Figure 8: Heaviest router-to-router link load and NoC area as the number of processors scales for different placement patterns. A tradeoff exists in placement pattern compactness and speed: runtime grows linearly and area grows superlinearly for the X-shaped placement pattern, and vice versa for the saturated square placement pattern.